OpenStudy (anonymous):

find the fourth term in the following sequence: A(n)= 5x (-3)n-1

OpenStudy (anonymous):

The fourth term is -135 because A(4)=5*(-3)^(4-1)=5*(-3)^3=5*-27=-135

OpenStudy (anonymous):

How does one know that there is an exponent, "^", involved here?

OpenStudy (anonymous):

Most problems that are given in textbooks use an exponent. Just an assumption.

OpenStudy (anonymous):

Well I am 60 years beyond graduating from High School and would never know that. One problem for me in dealing with problems on this site is that many of the problem expressions are ambiguous.